Okay, here it is:
We have 5mL of 80% acid. We want to dilute it with 2% acid to get 10% acid. How much 2% do we need? I tried to set it up, but can't seem to figure it out.
Thanks for any help you can give me.



Answer by stanbon(48558)

#1. Follow the acid.
Start with 80%(5)
Add 2%(x); x because the amount of fluid added is unknown
End up with: 10%(5+x)
EQUATION:
80%(5)+2%(x)=10%(5+x)
Multiply through by 100 to get:
80(5)+2x=10(5+x)
400+2x=50+10x
8x=350
x= 43.75ml (amount of 2% fluid you need to add)
Cheers,
Stan H.